Find the length of the chord of a circle of radius 100 feet subtended by an angle of 75◦ ?
1
Expert's answer
2018-03-01T12:32:08-0500
L=2пR/360*75
L=2*3.14*100/360*75=628/360*75=130.5 feet approximately